If business partner A invested $40,000 and business partner B invested $30,000 and they make a profit of $20,000, what would be business partner's B share?
Answer by rfer(16322)

4:3
4x+3x=20000
7x=20000
x=20000/7
x=2857.14
4x=11428.58 (A)
3x=8571.42 (B)